OpenStudy (anonymous):

Which expression is equivalent to the sine of the angle γ for the triangle of area A shown below? a. 2a/bc b. a/c c. b/c A d. Not enough information A. a B. b C. c D. d

OpenStudy (anonymous):

that a right triangle?

OpenStudy (anonymous):



OpenStudy (anonymous):

OK THIS ONE IS MORE ACCURATE :)

OpenStudy (anonymous):

Is that a right triangle?

OpenStudy (anonymous):

already asked that...

OpenStudy (anonymous):

No

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

see to solve this question draw a perpendicular on the side b. Now we know that the area of a right angled triangle is \[\frac{ 1 }{ 2 } \times base \times height\] Now we have two triangles with respective base Now use the area formula. so the area of the truangle A is the combined area of the two triangles. \[A= \frac{ 1 }{ 2 } h \times (b1+b2)\] Now sin(y) is h/c. Replace h from sin(y) using area eqn and you will get the answer as \[\frac{ 2A }{ bc }\]